OpenStudy (anonymous):

An impure Sample CaCO3 weighing 2.5g is completely reacted with 40ml of 0.5M H2So4.The Percentage of Purity of the Sample?

OpenStudy (callisto):

\[CaCO_3 + H_2SO_4 \rightarrow CaSO_4 + CO_2+H_2O\] 1. Find the no. of mole of H2SO4 2. No. of mole of CaCO3 reacted = no. of mole of H2SO4 3. Mass of CaCO3 reacted =No. of mole of CaCO3 reacted x molecular mass of CaCO3 4. % = answer in 3 / 2.5 x100% PS: I haven't done chemistry for long

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

\[H_2SO_4 + CaCO_3 --> CaSO_4 + H_2O + CO_2\] It is balanced, I hope this is what you are looking for. No. of moles = \[H_2SO_4=(1*2)+32+(4*16)= 2+32+64 = 98 mol\] Applying the law of conservation of mass = Mass of CaCO3 reacted =No. of mole of CaCO3 reacted x molecular mass of CaCO3

OpenStudy (anonymous):

Do you ge that @Yahoo!

OpenStudy (yrelhan4):

@rohangrr what does percentage purity actualy mean?